given:angle4 is an exterior angle of the triangle
Proves:m angle 4= angle 1+ angle 2
m angle 3+ m angle 4=180??
those that is Definition Supplementary angles???

Answer by vleith(2983)

Almost there.
Yes 3 and 4 are supplementary. So they total 180
The sum of the three angles in a triangle is always 180 too.
So 1 + 2 + 3 = 180
3 +4 = 180
1+2+3 = 3+4
1+2 = 4